# Butterfly in Spacetime: Inherent Instabilities in Stable Black Holes

### v2

@@ V Low-frequency instability
In above two examples, perturbed potentials $V_{p}$ themselves will trigger the instability if we turn off the Regge-Wheeler potential $V_{\text{RW}}$. Thus, roughly speaking, the instability can be understood as the competition between the destabilizing caused by perturbed potential $V_{p}$ and the stabilizing caused by BH. The following example will show that this is not the whole story. We specify the perturbed potential $V_{p}$ to be
@@
i.e, we consider a time-dependent perturbed potential that has Gaussian distribution in space but pseudo-random in time. We here focus on the situation $w\ll\mathcal{O}(1/r_{h})$.
Figure 3: Left panel: Time evolution of $\psi$ at $x=a$ in different situations with $w=0.02$ and $\sigma=4$. Right panel: Relationship between $\epsilon$ and $|a_{c}|$ for $w=0.001$ and $\sigma=4$. The red line is $\ln|a_{c}|=-\ln\epsilon-0.93$, which is obtained by fitting the most left 6 points.
@@
for sufficiently small $\epsilon$. Our numerical results suggest that, for the case $w\rightarrow 0^{+}$ the infinitesimal $\epsilon$ can still trigger the instability if the perturbed potential is sufficiently closed to the horizon. This effect appears unique to black holes as these same potentials do not destabilize Minkowski spacetime.
@@
Astrophysical BHs are inevitably surrounded by matter distributions, and even isolated BHs experience quantum fluctuations. These classical and quantum perturbations induce stochastic deformations in spacetime geometry. This study addresses a critical question: whether nominally stable BHs retain their stability when the geometry are deformed by such small perturbations. To model this effect, we map the deformations onto modifications of the effective potential governing s-wave scalar fields. Numerical analysis provides compelling evidence that BH stability can be disrupted by infinitesimal negative or stochastic modifications localized near the event horizon. Notably, certain perturbed potentials that stabilize systems in flat backgrounds can destabilize otherwise stable BHs. While our focus is on the event horizon, analogous behavior may occur near cosmological horizons in asymptotically de Sitter BHs. Our findings suggest that the stability of a stable BH is not robust against the tiny deformation of near-horizon geometry. Infinitesimal deformation of near-horizon geometry can turnover the stability of a stable BH. Due to the inevitable fluctuations of near horizon geometry, this intrinsic instability fundamentally constrains our ability to predict BH dynamics over extended timescales.
@@
As a toy model and very preliminary research, we here only considered static spherically symmetric case and numerically simulated the system in time domain. It would be interesting to consider Kerr black hole and study how spin would affect the phenomenon discovered here. Particularly, is there any relationship to turbulence of black holes discovered in Ref. Yang et al. (2015)? Could we use frequency analysis to verify the results and supply deeper understandings on the underlying mechanism?  [^1] The analysis of frequency domain for static negative bump and static stochastic potential will be addressed in an upcoming publication. Why does bump potential modulated by a time function in low frequency limit stabilize the flat spacetime but destabilize the black hole? More realistically, could we quantitatively estimate the disturbance intensity $\epsilon$ generated by the actual environment (such as accretion disks, dark matter halos) or quantum fluctuations, and whether the position $a$ satisfies the condition of $a<a_{c}$? We hope these questions could be addressed in the future through numerical simulations and further theoretical works.
@@
At the classical level, we emphasize that our results do not imply instability of the vacuum Schwarzschild solution, as no source exists to generate stochastic deformations on metric. For a BH surrounded by matter, our results implies that the system cannot keep static configuration for a long time in astrophysical settings, since uncontrollable stochastic matter fluctuations near the horizon will lead stochastic deformations on the geometry and inevitably trigger the instability identified here. The timescale that this instability becomes considerable depends on amplitude of the deformation. The probe scalar field studied here can act as a conduit, transferring energy from the environment to the BH and spatial infinity. Consequently, the final state evolves toward a Schwarzschild BH in vacuum, with increased horizon area and energy. However, if we posit an underlying quantum-gravitational framework beneath classical general relativity, our analysis indicates that even vacuum Schwarzschild BHs become unstable. This suggests that Schwarzschild BHs in a quantum-gravitational framework would gradually lose energy over time, potentially offering a new perspective on BH evaporation. Further investigation is needed to establish a direct connection and we hope some useful advances will be obtained in the future.
@@ Acknowledgements.
Appendix A Negative bump potential and dominate energy condition
In this appendix, we will give a concrete example that Regge-Wheeler potential has a local negative bump but dominate energy condition is satisfied.
@@ Appendix A Negative bump potential and dominate energy condition
We here assume that the event horizon locates at $x=-\infty$, where we have $r(-\infty)=r_{h}$ and $h(-\infty)=0$. The function $h(x)$ has asymptotically expression $h(x)=1-2M/x+\cdots$ with the total mass $M$ of the spacetime and $r^{\prime}(\infty)=1$. The Regge-Wheeler potential $V_{\mathrm{RW}}$ for s-wave scalar field then reads
@@
In static spherically symmetric case, the energy momentum tensor ${T^{\mu}}_{\nu}$ has a form ${T^{\mu}}_{\nu}=\text{diag}[-\rho(x),p_{x}(x),p_{T}(x),p_{T}(x)]$. The dominate energy condition then requires
@@
We now consider the near horizon region in general case. From Eqs. (17) and (18) one can find that there is no constraint on the values of $\rho-p_{x}$ and $p_{x}$. The values of $\rho$ and $p_{x}$ could be considerably large near horizon. Particularly, if an astrophysical BH is surrounded by sufficient matter localization around $x\sim a\ll-r_{h}$ and the dominate energy condition is satisfied, then it is possible
@@
as an example. To solve the Einstein equation, we note that the system is in vacuum and the solution is a Schwarzschild BH with Schwarzschild radius $r_{h}$ in the region $x<a-1/8$, of which the metric reads
$x=r_{h}+r_{h}\ln(1-r_{h}/r),~{}~{}r^{\prime}=h=1-r_{h}/r\,.$ (22)
We then integrate equations (17) and (18) into infinity with the initial values at $x=a-1/8$ given according to Eq. (22).
@@
In Fig. 4 we give a numerical result when we choose $a=-20,r_{h}=1$ and $b=0.2$. We see that there is negative bump around $x=a$. Thus, the negative bump potential can appear even if dominate energy condition is satisfied. We expect that, in general, localized spherical matter shells will introduce similar bumps in the Regge-Wheeler potential, possibly causing instabilities discussed in our main text.

### v3

@@ V Low-frequency instability
As we previously discussed, the stochastic local potential not only comes from the quantum fluctuations but also classical fluctuations. In Sec. IV, we focus on the quantum fluctuations with frequency near Planck scale, which will become $\mathcal{O}(1/r_{h})$ due to the extreme redshift near BH horizon. Since the proper frequency of classical fluctuations are much lower than the quantum fluctuations, we thus consider the low frequency $w\ll{\cal O}(1/r_{h})$ limit for the case of classical fluctuations near the BH horizons. We specify the perturbed potential $V_{p}$ to be
@@
i.e, we consider a time-dependent perturbed potential that has Gaussian distribution in space but pseudo-random in time. Here we set $w\ll 1$. In this situation, we will show that deformation potential itself will stabilize the flat background but can trigger the instability if it is closed to a BH.
Figure 3: Left panel: Time evolutions of $\psi$ at $x=a$ in different situations with $w=0.02$ and $\sigma=4$. Right panel: Relationship between $\epsilon$ and $|a_{c}|$ for $w=0.001$ and $\sigma=4$. The red line is $\ln|a_{c}|=-\ln\epsilon-0.93$, which is obtained by fitting the most left 6 points.
@@
for sufficiently small $\epsilon$. Our numerical results suggest that, for the case $w\rightarrow 0^{+}$ the infinitesimal $\epsilon$ can still trigger the instability if the perturbed potential is sufficiently close to the horizon. This effect appears unique to BHs as these same potentials do not destabilize Minkowski spacetime.
@@
Astrophysical BHs are inevitably surrounded by matter distributions, and even isolated BHs experience quantum fluctuations. These classical and quantum perturbations induce stochastic deformations in spacetime geometry. This study addresses a critical question: whether nominally stable BHs retain their stability when the geometry is deformed by such small perturbations. To model this effect, we map the deformations onto modifications of the effective potential governing s-wave scalar fields. Numerical analysis provides compelling evidence that BH stability can be disrupted by infinitesimal negative or stochastic modifications localized near the event horizon. Notably, certain perturbed potentials that stabilize systems in flat backgrounds can destabilize otherwise stable BHs. While our focus is on the event horizon, analogous behavior may occur near cosmological horizons in asymptotically de Sitter BHs. Our findings suggest that the stability of a stable BH is not robust against the tiny deformation of near-horizon geometry. Infinitesimal deformation of near-horizon geometry can overturn the stability of a stable BH. Due to the inevitable fluctuations of near horizon geometry, this intrinsic instability fundamentally constrains our ability to predict BH dynamics over extended timescales.
@@
As a toy model and very preliminary research, we here only considered static spherically symmetric case and numerically simulated the system in time domain. It would be interesting to consider Kerr BH and study how spin would affect the phenomenon discovered here. Particularly, is there any relationship to turbulence of BHs discovered in Ref. Yang et al. (2015)? Could we use frequency analysis to verify the results and supply deeper understandings on the underlying mechanism?  [^1] The analysis of frequency domain for static negative bump and static stochastic potential will be addressed in an upcoming publication. More realistically, could we quantitatively estimate the disturbance intensity $\epsilon$ generated by the actual environment (such as accretion disks, dark matter halos) or quantum fluctuations, and whether the position $a$ satisfies the condition of $a<a_{c}$? We hope these questions could be addressed in the future through numerical simulations and further theoretical work.
@@
At the classical level, we emphasize that our results do not imply instability of the vacuum Schwarzschild solution, as no source exists to generate stochastic deformations on metric. For a BH surrounded by matter, our results imply that the system cannot keep static configuration for a long time in astrophysical settings, since uncontrollable stochastic matter fluctuations near the horizon will lead to stochastic deformations on the geometry and inevitably trigger the instability identified here. The timescale that this instability becomes considerable depends on amplitude of the deformation. The probe scalar field studied here can act as a conduit, transferring energy from the environment to the BH and spatial infinity (See also Appendix. C). Consequently, the final state evolves toward a Schwarzschild BH in vacuum, with increased horizon area and energy. However, if we posit an underlying quantum-gravitational framework beneath classical general relativity, it becomes very interesting. We found that the quantum fluctuations could destabilize the Schwarzschild BH, implying that the scalar perturbations will exponentially grow with time and keep transferring energy to infinity. Will our finding be a potential mechanism for a new perspective on BH evaporation? This is a very interesting question. Further investigation is needed to establish a direct connection and we hope some useful advances will be obtained in the future.
@@ Acknowledgements.
Appendix A Negative bump potential and dominant energy condition
In this appendix, we will give a concrete example that Regge-Wheeler potential has a local negative bump but dominant energy condition is satisfied.
@@ Appendix A Negative bump potential and dominant energy condition
We here assume that the event horizon locates at $x=$-$\infty$, where we have $r(-\infty)=r_{h}$ and $h(-\infty)=0$. The function $h(x)$ has asymptotical expression $h(x)=1-2M/x+\cdots$ with the total mass $M$ of the spacetime and $r^{\prime}(\infty)=1$. The Regge-Wheeler potential $V_{\mathrm{RW}}$ for s-wave scalar field then reads
@@
In static spherically symmetric case, the energy momentum tensor ${T^{\mu}}_{\nu}$ has a form ${T^{\mu}}_{\nu}=\text{diag}[-\rho(x),p_{x}(x),p_{T}(x),p_{T}(x)]$. The dominant energy condition then requires
@@
We now consider the near horizon region in general case. From Eqs. (18) and (19) one can find that there is no constraint on the values of $\rho-p_{x}$ and $p_{x}$. The values of $\rho$ and $p_{x}$ could be considerably large near horizon. Particularly, if an astrophysical BH is surrounded by sufficient matter localization around $x\sim a\ll-r_{h}$ and the dominant energy condition is satisfied, then it is possible
@@
as an example. To solve the Einstein equation, we note that the system is in vacuum and the solution is a Schwarzschild BH with Schwarzschild radius $r_{h}$ in the region x<a1/8x<a-1/8, of which the metric reads
$x=r+r_{h}\ln(1-r_{h}/r),~~r^{\prime}=h=1-r_{h}/r\,.$ (23)
We then integrate equations (18) and (19) into infinity with the initial values at x=a1/8x=a-1/8 given according to Eq. (23).
@@
In Fig. 4 we give a numerical result when we choose $a=-20,r_{h}=1$ and $b=0.2$. We see that there is negative bump around $x=a$. Thus, the negative bump potential can appear even if dominant energy condition is satisfied. We expect that, in general, localized spherical matter shells will introduce similar bumps in the Regge-Wheeler potential, possibly causing instabilities discussed in our main text.
Appendix B Infinitesimal negative bump in static spherical star
In this appendix, we consider a static ideal star as background, then we find that infinitesimal negative bump cannot trigger the instability if it is placed sufficiently close to the star. Since an astrophysical star involves much complicated details of its interior, we here just us a toy model and only regard the star as a horizonless geometry with nonzero total mass. We consider a regular Hayward star. The metric of the Hayward star is  Hayward (2006)
$\text{d}s^{2}=-h(r)\text{d}t^{2}+\frac{\text{d}r^{2}}{h(r)}+r^{2}\text{d}\Omega^{2}\,,\quad h=1-\frac{2Mr^{2}}{r^{3}+2b^{2}M}\,.$ (24)
When $M<3\sqrt{3}/4b$, there is no horizon and the metric thus describe a regular star. For without the loss of generality, we set that $M=1,b=2$ as a concrete example. One important difference is that the tortoise coordinate $r_{*}$ here is defined as
$r_{*}=\int_{0}^{r}\frac{\text{d}x}{h(x)}\in[0,\infty)$ (25)
This leads fundamental changes: For BH, “infinitely close to horizon” means $r_{*}\rightarrow$-$\infty$; For a star, “infinitely close to star” means $r_{*}$ is still finite and can at most decrease to zero. This difference is universal for general spherically symmetric stars and BHs.
Figure 5: Time evolutions of $\psi$ at $x=a$ in different situations with $\epsilon=0.01$ and 0.001
In Fig. 5, we numerically shows two cases of $a=0$: $\epsilon=0.01$ and $0.001$. We choose $a=0$ since this stands for the case that is closest to the star. Clearly, sufficiently small negative bump cannot trigger the instability when it is “sufficiently close to the star”. [^2] When the negative bump is sufficiently far way from the star, the instability can still be triggered but this is the effect of negative bump in flat space.
Appendix C Energy flow of the scalar field
To explain how the probe scalar field would serve as the conduit, we can compute the energy fluxes towards event horizon and infinity. For a light massless scalar field, the energy momentum tensor
$T^{\mu\nu}=\nabla^{\mu}\Psi\nabla^{\nu}\Psi-\frac{1}{2}g^{\mu\nu}(\nabla\Psi)^{2}\,$, (26)
where $g^{\mu\nu}$ denotes the metric of Schwarzschild black hole background and $\Psi=\psi/r$. It needs to note that asymptotical flatness requires $\Psi$ to be zero at infinity but $\psi$ could be finite. Since the background spacetime is assumed to be static with timelike Killing vector $\xi^{\mu}=(\partial/\partial t)^{\mu}$, then we can define the an energy current $J^{\mu}=-T^{\mu\nu}\xi_{\nu}$ and it satisfies $\nabla_{\mu}J^{\mu}=0$.
In order to study the energy fluxes at horizon and infinity, let us consider a family of spacelike hypersurfaces $\{\Sigma_{s}\}$ outside event horizon, which gives a foliation on the spacetime region between $r_{h}$ and $r=r_{0}\gg r_{h}$. We assume that this foliation preserve the spherical symmetry. The boundaries of every $\Sigma_{s}$ are given by $r=r_{h}$ and $r=r_{0}$. The energy in one hypersurface $\Sigma_{s}$ is given by
$E_{s}=\int_{\Sigma_{s}}J^{\mu}\text{d}\Sigma_{\mu}\,.$ (27)
Here $\text{d}\Sigma_{\mu}$ is the future-toward directed volume element. Let us consider two different hypersurface $\Sigma_{1}$ and $\Sigma_{2}$, see Fig. 6.
Figure 6: Penrose diagram on the exterior of horizon and a family of spacelike hypersurfaces $\{\Sigma_{s}\}$. Here $\mathcal{N}$ is boundary belong to horizon and $S$ is the boundary at $r=r_{0}\gg r_{h}$. In the limit $r_{0}\rightarrow\infty$, we have $S\subset i_{0}$. It should be note that $\xi^{\mu}$ is ‘outward directed’ normal to $\mathcal{N}$, as determined by continuity.
The difference of energies between these two hypersurfaces is
$\begin{split}&E_{2}-E_{1}=\int_{\Sigma_{2}}J^{\mu}\text{d}\Sigma_{\mu}-\int_{\Sigma_{1}}J^{\mu}\text{d}\Sigma_{\mu}\\ &=\oint_{\partial V}J^{\mu}\text{d}\Sigma_{\mu}-\left(\int_{\mathcal{N}}J^{\mu}\text{d}\Sigma_{\mu}+\int_{S}J^{\mu}\text{d}\Sigma_{\mu}\right)\,.\end{split}$ (28)
Here $\mathcal{N}$ is boundary belong to horizon, $S$ is the boundary at $r=r_{0}$ and $V$ is the spacetime region surrounded by $\Sigma_{1},\Sigma_{2},\mathcal{N}$ and $S$. The hypersurface $S$ is define by $r=r_{0}$ and $t\in[t_{1},t_{2}]$. The null hypersurface $\mathcal{N}$ is defined by $r=r_{h}$ and $v\in[v_{1},v_{2}]$, where $(\partial/\partial v)^{\mu}$ is the future-directed generator of event horizon. At event horizon, $\xi^{\mu}$ is both normal and tangent to event horizon and we take $(\partial/\partial v)^{\mu}=\xi^{\mu}$.
After using Gaussian theorem and $\nabla_{\mu}J^{\mu}=0$, we then have
$E_{2}-E_{1}=-\left(\int_{\mathcal{N}}J^{\mu}\text{d}\Sigma_{\mu}+\int_{S}J^{\mu}\text{d}\Sigma_{\mu}\right)\,.$ (29)
The above two integrations describe the energy fluxes at horizon and $r=r_{0}$, respectively. Since $r=r_{0}$ is a timelike surface and its unit outer normal vector reads $n^{\mu}=(\partial/\partial r)^{\mu}$ (we assume that $r_{0}\gg r_{h}$). The energy flux at $r=r_{0}$ then is given by
$\begin{split}\int_{S}J^{\mu}\text{d}\Sigma_{\mu}&=-4\pi r_{0}^{2}\int_{t_{1}}^{t_{2}}\text{d}t\dot{\Psi}\partial_{r}\Psi|_{r=r_{0}}\\ &=\left.-4\pi\int_{t_{1}}^{t_{2}}\text{d}t\dot{\psi}\left(\frac{\partial\psi}{\partial r}-r^{-1}\psi\right)\right|_{r=r_{0}}\,.\end{split}$ (30)
The “no-reflect boundary condition” at the infinity shows that
$\left.\dot{\psi}+\frac{\partial\psi}{\partial r}\right|_{r\rightarrow\infty}=0\,$,
and asymptotical flatness requires $\psi/r=\Psi\rightarrow 0$. The energy flux at $S$ then is positive and is toward infinity. Since $\mathcal{N}$ is a null surface and its outer normal vector reads $\xi^{\mu}$ (we assume that $r_{0}\gg r_{h}$). The energy flux at $\mathcal{N}$ then is given by
$\begin{split}&\int_{\mathcal{N}}J^{\mu}\text{d}\Sigma_{\mu}=-\int_{\mathcal{N}}J^{\mu}\xi_{\mu}\text{d}\Sigma\\ &=4\pi r_{h}^{2}\int_{v_{1}}^{v_{2}}\text{d}t\dot{\Psi}\partial_{t}\Psi|_{r=r_{0}}=4\pi\int_{v_{1}}^{v_{2}}\text{d}v\dot{\psi}^{2}\,.\end{split}$ (31)
Here we have used the fact that $\xi^{\mu}$ is null vector at horizon. The energy flux towards interior of event horizon then is also positive.
In the stable case, any fluctuation of test scalar field will decay to zero fast, so energy fluxes at $\mathcal{N}$ and $S$ will also decay to zero fast. However, in the situations discussed in main text, the fluctuation of test scalar field will exponentially increase, so energy fluxes at $\mathcal{N}$ and $S$ will be more and more. Therefore the scalar field acts as a conduit, transferring energy from the environment into the black hole and out to infinity.